Im having a problem understanding how to slove a problem when the varible is next to 3 numbers: 5x+3=2x-(5-3x) i dont know what to do with the 5x or where to use the distributive property.


i have a problem with understanding where to use the distributive property and where i should cancell out numbers in this equation:2(3n+6)=3(12-2n)

thanks for the help:)

Answer by ankor@dixie-net.com(22740)

5x + 3 = 2x -(5-3x)
the neg in front of the brackets means
5x + 3 = 2x -1(5-3x)
Distribute the -1 in front of the brackets means you multiply what's inside brackets by -1
5x + 3 = 2x - 5 + 3x
Collect like terms on the right
5x + 3 = 2x + 3x - 5
5x + 3 = 5x - 5
subtract 3 from both sides:
5x = 5x - 5 - 3
5x = 5x - 8
obviously this is not a valid equation, we can say x is undefined
:
:
2(3n+6) = 3(12-2n)
Distribute the 2 and the 3, multiply what's inside the brackets
6n + 12 = 36 - 6n
Add 6n to both sides
6n + 6n + 12 = 36
subtract 12 from both sides, add the n's
12n = 36 - 12
12n = 24
divide both sides by 12
n = 24/12
n = 2
:
Check this in the original problem
2(3n+6) = 3(12-2n)
replace n with 2
2(3(2)+6) = 3(12-2(2))
2(6+6) = 3(12-4)
2(12) = 3(8)
24 = 24, confirms our solution of n=2